Max has $3.00 in change in his pocket, consisting of quarters, dimes and nickels. The number of
quarters is one-half the number of nickels and the number of dimes is three more than the number
of quarters. How many nickels, dimes and quarters does Max have in his pocket?

Answer by VFBundy(438)

Number of quarters = Q
Number of nickels = 2Q
Number of dimes = Q + 3

0.25Q + 0.05(2Q) + 0.10(Q + 3) = 3.00

0.25Q + 0.10Q + 0.10Q + 0.30 = 3.00

0.45Q + 0.30 = 3.00

0.45Q = 2.70

Q = 6

Number of quarters = Q = 6
Number of nickels = 2Q = 12
Number of dimes = Q + 3 = 9